# Fluctuation induced piezomagnetism in local moment altermagnets

### v1

@@ III Conclusions
Here we have identified a piezomagnetic response that instead grows with temperature, because it is driven by thermally excited magnons and is described by formula (4). This piezomagnetic response is due to transversal magnetic fluctuations and can thus also be expected for systems with fixed local moments. It reaches a maximum in a temperature region just below $T_{\textsc{n}}$. Interestingly, the thermally-induced piezomagnetism is dominant for materials with small magnetic moments $\mu_{s}\ll\mu_{\textsc{b}}$ in the temperature regime $T\lesssim T_{\textsc{n}}$. We have also shown that in presence of magnetic fluctuations a spin current is generated by an applied temperature gradient due to different magnon branches carrying opposite magnetic moment. The spin carried by the heat current couples to the direction of that current. These results are easily generalized to higher dimensions and different lattice geometries.
@@ Acknowledgments
We thank Jorge Facio and Oleg Jansson for fruitful discussions. This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the Sonderforschungsbereich SFB 1143, grant No. YE 232/1-1, and under Germany’s Excellence Strategy through the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter – ct.qmat (EXC 2147, project-ids 390858490 and 392019).
@@ Appendix A Dispersion relation for the Heisenberg checkerboard model
$\begin{split}&\mathcal{H}^{(2)}=\sum\limits_{\bm{k}\in 1.\text{BZ}}\biggl{[}4J\cos\frac{a_{0}k_{x}}{2}\cos\frac{a_{0}k_{y}}{2}\hat{m}_{1\alpha}(\bm{k})\hat{m}_{2\alpha}(-\bm{k})\\ &+\left(2J-2\tilde{J}_{y}\sin^{2}\frac{k_{y}a_{0}}{2}+K+\frac{\mu_{s}B}{2}\right)\hat{m}_{1\alpha}(\bm{k})\hat{m}_{1\alpha}(-\bm{k})\\ &+\left(2J-2\tilde{J}_{x}\sin^{2}\frac{k_{x}a_{0}}{2}+K-\frac{\mu_{s}B}{2}\right)\hat{m}_{2\alpha}(\bm{k})\hat{m}_{2\alpha}(-\bm{k})\biggr{]},\end{split}$ (15)
where the summation over the repeating index $\alpha\in\{x,y\}$ is assumed. With (15), we write Eqs. (14) in the form
@@
$\mathbb{M}=\begin{bmatrix}0&-(\mathcal{B}_{y}+b)&0&-\mathcal{A}\\ \mathcal{B}_{y}+b&0&\mathcal{A}&0\\ 0&\mathcal{A}&0&\mathcal{B}_{x}-b\\ -\mathcal{A}&0&-(\mathcal{B}_{x}-b)&0\end{bmatrix}$ (17a)
with $b=\mu_{s}B/(4J)$ and
$\mathcal{A}=\cos\frac{k_{x}a_{0}}{2}\cos\frac{k_{x}a_{0}}{2},$ (17b)
$\mathcal{B}_{\alpha}=1+\frac{\kappa}{2}-\epsilon_{\alpha}\sin^{2}\frac{k_{\alpha}a_{0}}{2}.$ (17c)
Parameters $\omega_{0}$, $\kappa$, and $\epsilon_{\alpha}$ are defined in the main text. System (16) has solution $\bm{\xi}=\bm{\xi}_{0}e^{-i\omega t}$, which is nontrivial ($\bm{\xi}_{0}\neq\bm{0}$) if $\omega=i\omega_{0}\lambda_{\nu}$ with $\lambda_{\nu}$ being the eigenvalues of matrix $\mathbb{M}$. The eigenvalues $\lambda_{\nu}$ are imaginary and compose two complex-conjugated pairs. The pair of the non-negative eigenfrequencies is presented in (2). Note that $\gamma B=\omega_{0}b$.
Appendix B Magnetization for low temperature
The explicit form of function $f(\epsilon_{x},\epsilon_{y})$ is as follows
$\begin{split}&f=\frac{2}{\pi^{2}\delta\epsilon}\int\limits_{0}^{\pi}\biggl{\{}\left[1+\left(1+\frac{\kappa}{2}\right)(\bar{\epsilon}+\delta\epsilon\cos\chi)\right]^{2}\\ &-\Delta^{2}(\delta\epsilon+\bar{\epsilon}\cos\chi)^{2}\biggr{\}}^{-1}(\delta\epsilon+\bar{\epsilon}\cos\chi)\,\mathrm{d}\chi,\end{split}$ (18)
where we introduced $\delta\epsilon=(\epsilon_{x}-\epsilon_{y})/2$ and $\bar{\epsilon}=(\epsilon_{x}+\epsilon_{y})/2$. In the limit $\delta\epsilon\ll 1$, we obtain
$f=\frac{2/\pi}{\left\{\left[1+(1+\frac{\kappa}{2})\bar{\epsilon}\right]^{2}-\Delta^{2}\bar{\epsilon}^{2}\right\}^{3/2}}+\mathcal{O}(\delta\epsilon^{2}).$ (19)
Assuming $\bar{\epsilon}\ll 1$, one obtains the approximation presented in the main text.
Appendix C Tensor of thermal spin conductivity
The exact expressions for $\sigma_{\alpha\beta}$ defined in (7) are as follows
and
$\begin{split}&\sigma_{yy}=-\gamma k_{\textsc{b}}\beta^{\prime 2}\frac{\tau_{\text{rlx}}\omega_{0}}{64}\iint\limits_{-\pi}^{\pi}\frac{\mathrm{d}q_{x}\mathrm{d}q_{y}}{(2\pi)^{2}}\sin^{2}q_{y}\\ &\times\Biggl{\{}\frac{(F_{q}+\Omega_{q}^{-})\left[\frac{1}{F_{q}}\left(\cos^{2}\frac{q_{x}}{2}-\epsilon_{y}\left(1+\frac{\kappa}{2}-\Omega_{q}^{+}\right)\right)-\epsilon_{y}\right]^{2}}{\sinh^{2}\left[\frac{\beta^{\prime}}{2}(F_{q}+\Omega_{q}^{-})\right]}\\ &-\frac{(F_{q}-\Omega_{q}^{-})\left[\frac{1}{F_{q}}\left(\cos^{2}\frac{q_{x}}{2}-\epsilon_{y}\left(1+\frac{\kappa}{2}-\Omega_{q}^{+}\right)\right)+\epsilon_{y}\right]^{2}}{\sinh^{2}\left[\frac{\beta^{\prime}}{2}(F_{q}-\Omega_{q}^{-})\right]}\Biggr{\}}\end{split}$ (21)
and $\sigma_{xy}=\sigma_{yx}=0$.
Appendix D Numerical simulations of the Heisenberg checkerboard
@@
D.1 Simulation of spinwaves
@@
To simulate spinwaves we considered a system with a size of $N_{1}\times N_{2}=500\times 500$ magnetic moments. The simulations are carried out in two steps. In the first step, we simulate the dynamics of the system in the external magnetic field $\bm{B}_{i}=B_{0}\,\text{sinc}\left(2\pi\ \bm{k}\cdot\bm{R}_{i}\right)\text{sinc}\left[2\pi\omega_{0}\left(t-t_{0}\right)\right]$, where $t_{0}=t_{\text{sim}}/2$ is a center of temporal part of field profile and $H_{0}$ is an amplitude of the applied field. The simulations are performed in the low damping regime with $\alpha=10^{-3}$.
@@
In the second step we performed a space-time transform for the complex-valued parameter $m_{i}^{\textsc{x}}(t)+i\ m_{i}^{\textsc{y}}(t)$. The resulting eigenfrequencies are plotted in Fig. 2.
@@
Figure 5: Temperature dependence of perpendicular component of Néel vector $n_{z}$ is shown for $\epsilon_{x}=-\epsilon_{y}=0.1$ and for two different anisotropy values. Symbols correspond to the results of numerical simulations.
D.2 Simulation of temperature-induced effects
@@
Here we consider a system with a size of $N_{1}\times N_{2}=1000\times 1000$ magnetic moments. In the simulations, we consider the temperature in the range $T\in\left[0.15;\,0.55\right]\hbar\omega_{0}/k_{\textsc{b}}$. The simulations were performed for a low damping regime with $\alpha=10^{-3}$ for a long time scale with $\alpha t\omega_{0}\gg 1$. The averaged perpendicular net magnetization and Néel vectors are presented in Figs. 3 and 5, respectively. Note that the averaged in-plane components of the net magnetization vanish.

### v2

@@ III Conclusions
Here we have identified a piezomagnetic response that instead grows with temperature, because it is driven by thermally excited magnons and is described by formula (5). This piezomagnetic response is due to transversal magnetic fluctuations and can thus also be expected for systems with fixed local moments. It reaches a maximum in a temperature region just below $T_{\textsc{n}}$. Interestingly, the thermally-induced piezomagnetism is dominant for materials with small magnetic moments $\mu_{s}\ll\mu_{\textsc{b}}$ in the temperature regime $T\lesssim T_{\textsc{n}}$. We have also shown that in presence of magnetic fluctuations a spin current is generated by an applied temperature gradient due to different magnon branches carrying opposite magnetic moment. The spin carried by the heat current couples to the direction of that current. These results are easily generalized to higher dimensions and different lattice geometries. For instance, we show that the model of an altermagnet with a structure of rutile results in the same low-temperature behavior of $\eta$ and $\sigma$ as the checkerboard model.
@@
A way to observe the fluctuation-induced piezomagnetism may be to apply the experimental approach of Ref. 10 to a $d$-wave altermagnet, as for the originally considered $g$-wave altermagnet, the effect is averaged out because of the higher symmetry of the magnon dispersion.
@@ Acknowledgments
We thank Jorge Facio and Oleg Jansson for fruitful discussions and Ulrike Nitzsche for technical support. This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the Sonderforschungsbereich SFB 1143, grant No. YE 232/2-1, and under Germany’s Excellence Strategy through the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter – ct.qmat (EXC 2147, project-ids 390858490 and 392019).
@@ Appendix A Dispersion relation for the Heisenberg checkerboard model
$\begin{split}&\mathcal{H}^{(2)}=4J\sum\limits_{\bm{k}\in 1.\text{BZ}}\biggl{[}\mathcal{A}_{\bm{k}}\hat{m}_{1\alpha}(\bm{k})\hat{m}_{2\alpha}(-\bm{k})\\ &+\frac{\mathcal{B}_{\bm{k}}+\mathcal{C}_{\bm{k}}}{2}\hat{m}_{1\alpha}(\bm{k})\hat{m}_{1\alpha}(-\bm{k})+\frac{\mathcal{B}_{\bm{k}}-\mathcal{C}_{\bm{k}}}{2}\hat{m}_{2\alpha}(\bm{k})\hat{m}_{2\alpha}(-\bm{k})\biggr{]},\end{split}$ (19)
where the summation over the repeating index $\alpha\in\{x,y\}$ is assumed, $\mathcal{A}_{\bm{k}}=\cos\frac{k_{x}a_{0}}{2}\cos\frac{k_{y}a_{0}}{2}$, $\mathcal{B}_{\bm{k}}=1+\frac{\kappa}{2}-\Omega_{\bm{k}}^{+}$, $\mathcal{C}_{\bm{k}}=\Omega_{\bm{k}}^{-}+b$ with $b=\mu_{s}B/(4J)$ Parameters $\kappa$, $\epsilon_{\alpha}$ as well as $\Omega_{\bm{k}}^{\pm}$ are defined in the main text. With (19), we write Eqs. (18) in the form
@@
$\mathbb{M}=\begin{bmatrix}0&-(\mathcal{B}_{\bm{k}}+\mathcal{C}_{\bm{k}})&0&-\mathcal{A}_{\bm{k}}\\ \mathcal{B}_{\bm{k}}+\mathcal{C}_{\bm{k}}&0&\mathcal{A}_{\bm{k}}&0\\ 0&\mathcal{A}_{\bm{k}}&0&\mathcal{B}_{\bm{k}}-\mathcal{C}_{\bm{k}}\\ -\mathcal{A}_{\bm{k}}&0&-(\mathcal{B}_{\bm{k}}-\mathcal{C}_{\bm{k}})&0\end{bmatrix}$ (21)
System (20) has solution $\bm{\xi}=\bm{\xi}_{0}e^{-i\omega t}$, which is nontrivial ($\bm{\xi}_{0}\neq\bm{0}$) if $\omega=i\omega_{0}\lambda_{\nu}$ with $\lambda_{\nu}$ being the eigenvalues of matrix $\mathbb{M}$. The eigenvalues $\lambda_{\nu}$ are imaginary and compose two complex-conjugated pairs. The pair of the non-negative eigenfrequencies $\omega_{1,2}=\omega_{0}(\sqrt{\mathcal{B}_{\bm{k}}^{2}-\mathcal{A}_{\bm{k}}^{2}}\pm\mathcal{C}_{\bm{k}})$ coincides with (2). Note that $\gamma B=\omega_{0}b$.
Appendix B Numerical simulations of the Heisenberg checkerboard
Figure 6: (a) – Time step dependence of the perpendicular Néel vector component for different reduced temperatures for Heun’s integration scheme. (b) – Comparison between the dynamics of the perpendicular Néel vector component $\overline{n}_{z}(t)$ with different system size. (c) – Mean energy per spin of the 1D Heisenberg chain, as a function of temperature. Symbols show the result of numerical simulations and the solid line corresponds to the analytical prediction. In all cases simulations were performed for the Gilbert damping parameter $\alpha=0.1$.
@@ Appendix B Numerical simulations of the Heisenberg checkerboard
B.1 Convergence of numerical results
Finite time step. For micromagnetic simulations at zero temperature, the minimum time step is a well defined quantity since the largest field (usually the exchange field) essentially defines the precession frequency. However, for spin-lattice simulations using the stochastic Landau–Lifshitz equation, the effective field becomes temperature dependent, see Eqs. (22) and (23). The consequence of this is that for spin-lattice models the most difficult region to integrate is near the Néel temperature. Errors in the integration of the system will be apparent from a non-converged value for the average Néel vector. This gives a relatively simple case which can then be used to test the stability of integration schemes for the stochastic Landau–Lifshitz equations. A plot of the mean perpendicular component of Néel vector $n_{z}$ as a function of temperature is shown in Fig. 6(a) for a system consisting of $500\times 500$ magnetic moments, where we take into account only nearest-neighbors AFM exchange, i.e. $\tilde{J}_{x}=\tilde{J}_{y}=0$, and easy-axial anisotropy $\kappa=0.5$.
The data in Fig. 6(a) shows that for low temperatures reasonably large time steps of $\omega_{0}\delta t=10^{-1}$  [^3] For material parameters of RuO2 [13] the corresponding characteristic time scale parameter is $1/\omega_{0}\approx 45$ fs. give the correct solution of the Landau–Lifshitz equations, while near the Néel temperature the deviations from the correct equilibrium value are significant. Consequently for higher temperatures time steps of $\omega_{0}\delta t=10^{-3}-10^{-2}$ are necessary.
Finite size. In the next step we check the system size dependence of our results for samples with nearest-neighbors AFM exchange, i.e. $\tilde{J}_{x}=\tilde{J}_{y}=0$, and easy-axial anisotropy $\kappa=0.5$. A plot of the mean perpendicular component of Néel vector $n_{z}$ as a function of sample size is shown in Fig. 6(b) for temperature $T=0.1\hbar\omega_{0}/k_{\textsc{b}}$. The data in Fig. 6(b) shows that a small sample ($100\times 100$) gives almost the same value of $\overline{n}_{z}$ as a large sample ($1000\times 1000$). In the following, we consider samples with a size of $500\times 500$ magnetic moments.
1D Heisenberg chain. The simplest model of classical interacting spins is the 1D Heisenberg chain with nearest-neighbor interactions. For this system, an analytical expression for the mean energy per spin is known [28] $\overline{H}/J=k_{\textsc{b}}T/J-1/\tanh\left[J/\left(k_{\textsc{b}}T\right)\right]$. The comparison of the Heun’s integration method with the analytical prediction as function of temperature is shown in Fig. 6(c). Here, we consider 1D chain with a length of 500 magnetic moments with nearest-neighbors ferromagnetic exchange and zero anisotropy $\kappa=0$.
B.2 Simulation of spinwaves
@@
To simulate spinwaves we considered a system with a size of $N_{1}\times N_{2}=500\times 500$ magnetic moments. The simulations are carried out in two steps. In the first step, we simulate the dynamics of the system in the external magnetic field $\bm{B}_{i}=B_{0}\,\text{sinc}\left(2\pi\omega_{0}\ \bm{k}\cdot\bm{R}_{i}\right)\text{sinc}\left[2\pi\left(t-t_{0}\right)\right]$, where $t_{0}=t_{\text{sim}}/2$ is a center of temporal part of field profile and $H_{0}$ is an amplitude of the applied field. The simulations are performed in the low damping regime with $\alpha=10^{-3}$.
@@
In the second step, we performed a space-time transform for the complex-valued parameter $m_{i}^{\textsc{x}}(t)+i\ m_{i}^{\textsc{y}}(t)$. The resulting eigenfrequencies are plotted in Fig. 2.
@@
Figure 7: Temperature dependence of perpendicular component of Néel vector $n_{z}$ is shown for $\epsilon_{x}=-0.4$, $\epsilon_{y}=-0.2$, and for two different anisotropy values. Symbols correspond to the results of numerical simulations. Dashed lines show fitting $\bar{n}_{z}=\left(1-T/T_{\textsc{n}}\right)^{\alpha}$ with (i) $T_{\textsc{n}}\approx 0.33\hbar\omega_{0}/k_{\textsc{b}}$ and $\alpha\approx 0.29$ for $\kappa=0.1$, and (ii) $T_{\textsc{n}}\approx 0.41\hbar\omega_{0}/k_{\textsc{b}}$ and $\alpha=0.27$ for $\kappa=0.5$.
B.3 Simulation of temperature-induced effects
@@
Here we consider a system with a size of $N_{1}\times N_{2}=500\times 500$ magnetic moments. In the simulations, we consider the temperature in the range $T\in\left[0.05;\,0.45\right]\hbar\omega_{0}/k_{\textsc{b}}$. The simulations were performed for a low damping regime with $\alpha=10^{-3}$ for a long time scale with $\alpha t\omega_{0}\gg 1$. The averaged perpendicular net magnetization and Néel vectors are presented in Figs. 4 and 7, respectively. Note that the averaged in-plane components of the net magnetization vanish.
Appendix C Fluctuation-induced piezomagnetism and thermal spin conductivity for rutiles
Here we consider a modified model of a rutile bilayer film previously proposed in [13], see Fig 3. Magnetic Hamiltonian consists of several contributions: $\mathcal{H}=\mathcal{H}_{\textsc{afm}}+\mathcal{H}_{\textsc{fm}}+\mathcal{H}_{\textsc{alt}}+\mathcal{H}_{\textsc{a}}+\mathcal{H}_{\textsc{z}}$. The leading one is the antiferromagnetic exchange between sublattices
$\mathcal{H}_{\textsc{afm}}=J\sum\limits_{\bm{R}_{\bm{n}}}\sum\limits_{\varsigma\in\Xi_{\times}}\bm{m}_{1}(\bm{R}_{\bm{n}})\cdot\bm{m}_{2}(\bm{R}_{\bm{n}}+\bm{\delta R}_{\varsigma}),$ (24)
where $J>0$ and $\Xi_{\times}=\{\nearrow,\nwarrow,\swarrow,\searrow\}$ and we introduced the displacement vectors $\bm{\delta R}_{\nearrow}=-\bm{\delta R}_{\swarrow}=\frac{a_{0}}{2}(\bm{e}_{1}+\bm{e}_{2})$, $\bm{\delta R}_{\nwarrow}=-\bm{\delta R}_{\searrow}=\frac{a_{0}}{2}(-\bm{e}_{1}+\bm{e}_{2})$. Additionally we take into account the nearest-neighbours ferromagnetic exchange interaction within each of the sublattices
$\begin{split}\mathcal{H}_{\textsc{fm}}=&-\frac{J_{\textsc{fm}}}{2}\sum\limits_{\bm{R}_{\bm{n}}}\sum\limits_{\varsigma\in\Xi_{+}}\bm{m}_{1}(\bm{R}_{\bm{n}})\cdot\bm{m}_{1}(\bm{R}_{\bm{n}}+\bm{\delta R}_{\varsigma})\\ &-\frac{J_{\textsc{fm}}}{2}\sum\limits_{\bm{R}^{\prime}_{\bm{n}}}\sum\limits_{\varsigma\in\Xi_{+}}\bm{m}_{2}(\bm{R}^{\prime}_{\bm{n}})\cdot\bm{m}_{2}(\bm{R}^{\prime}_{\bm{n}}+\bm{\delta R}_{\varsigma}),\end{split}$ (25)
where $J_{\textsc{fm}}>0$ and $\Xi_{+}=\{\uparrow,\downarrow,\leftarrow,\rightarrow\}$ and the displacement vectors are $\bm{\delta R}_{\rightarrow}=-\bm{\delta R}_{\leftarrow}=a_{0}\bm{e}_{x}$ and $\bm{\delta R}_{\uparrow}=-\bm{\delta R}_{\downarrow}=a_{0}\bm{e}_{y}$. The altermagnetic properties are modeled by the next-nearest exchange couplings acting within each of the sublattices
$\begin{split}\mathcal{H}_{\textsc{alt}}=&\frac{\tilde{J}_{1}}{2}\sum\limits_{\bm{R}_{\bm{n}}}\bm{m}_{1}(\bm{R}_{\bm{n}})\cdot\bigl{[}\bm{m}_{1}(\bm{R}_{\bm{n}}+2\bm{\delta R}_{\searrow})+\bm{m}_{1}(\bm{R}_{\bm{n}}+2\bm{\delta R}_{\nwarrow})-\bm{m}_{1}(\bm{R}_{\bm{n}}+2\bm{\delta R}_{\swarrow})-\bm{m}_{1}(\bm{R}_{\bm{n}}+2\bm{\delta R}_{\nearrow})\bigr{]}\\ +&\frac{\tilde{J}_{2}}{2}\sum\limits_{\bm{R}^{\prime}_{\bm{n}}}\bm{m}_{2}(\bm{R}^{\prime}_{\bm{n}})\cdot\bigl{[}-\bm{m}_{2}(\bm{R}^{\prime}_{\bm{n}}+2\bm{\delta R}_{\searrow})-\bm{m}_{2}(\bm{R}^{\prime}_{\bm{n}}+2\bm{\delta R}_{\nwarrow})+\bm{m}_{2}(\bm{R}^{\prime}_{\bm{n}}+2\bm{\delta R}_{\swarrow})+\bm{m}_{2}(\bm{R}^{\prime}_{\bm{n}}+2\bm{\delta R}_{\nearrow})\bigr{]}\end{split}$ (26)
In contrast to the previously considered model of a rutile film [13], here we assume that generally $\tilde{J}_{1}\neq\tilde{J}_{2}$. The latter model the action of a stress applied along diagonal directions [110] or [$\bar{1}$10]. We assume that $|\tilde{J}_{1,2}|\ll J+J_{\textsc{fm}}$, thus the ground state is a collinear-compensated one. We also take into account the perpendicular easy-axial anisotropy
$\mathcal{H}_{\textsc{a}}=-K\left[\sum\limits_{\bm{R}_{\bm{n}}}m_{1z}^{2}(\bm{R}_{\bm{n}})+\sum\limits_{\bm{R}^{\prime}_{\bm{n}}}m_{2z}^{2}(\bm{R}^{\prime}_{\bm{n}})\right]$ (27)
with $K>0$ and Zeeman interaction with magnetic field $\bm{B}=B\bm{e}_{z}$
$\mathcal{H}_{\textsc{z}}=-\mu_{s}B\left[\sum\limits_{\bm{R}_{\bm{n}}}m_{1z}(\bm{R}_{\bm{n}})+\sum\limits_{\bm{R}^{\prime}_{\bm{n}}}m_{2z}(\bm{R}^{\prime}_{\bm{n}})\right]$ (28)
In reciprocal space, Hamiltonian of the spin waves excited over the ground state $\bm{m}_{1}=\bm{e}_{z}$, $\bm{m}_{2}=-\bm{e}_{z}$ has the structure of (19) with $\mathcal{A}_{\bm{k}}=\cos\frac{k_{x}a_{0}}{2}\cos\frac{k_{y}a_{0}}{2}$, $\mathcal{B}_{\bm{k}}=1+\frac{\kappa}{2}+\upsilon(\sin^{2}\frac{k_{x}a_{0}}{2}+\sin^{2}\frac{k_{y}a_{0}}{2})+\delta\epsilon\sin(k_{x}a_{0})\sin(k_{y}a_{0})$, $\mathcal{C}_{\bm{k}}=\bar{\epsilon}\sin(k_{x}a_{0})\sin(k_{y}a_{0})+b$. Here $\kappa$ and $b$ are the same as previously, $\upsilon$, $\bar{\epsilon}$, $\delta\epsilon$ are defined in the main text. The corresponding dispersion relation has the form of (2a) with $F_{\bm{k}}=\sqrt{\mathcal{B}_{\bm{k}}^{2}-\mathcal{A}_{\bm{k}}^{2}}$ and $\Omega_{\bm{k}}^{-}=\mathcal{C}_{\bm{k}}$. The derivation is analogous to one presented in Appendix A.
In the low temperature limit, the corresponding asymptotic for $\eta$ coincides with (31a) up to the value of the function $f=-16\bar{\epsilon}(1+\frac{\kappa}{2})/[1+\upsilon(2+\kappa)]^{3}$. Here we also assumed that $\bar{\epsilon}\sqrt{\kappa}\ll 1$. For the parameters of RuO2 listed in supplemental materials of Ref. 13 we estimate $\eta\approx 0.13\mu_{\textsc{b}}/\mathrm{nm}^{2}$ for the room temperature.
Let us consider the thermal spin conductivity for the rutile model. We limit ourselves to the case $\delta\epsilon=0$, $b=0$, and we also denote $\epsilon=\bar{\epsilon}$. Taking into account the symmetry properties of $\Omega_{\bm{k}}^{-}$ and $F_{\bm{k}}$ in this case, we obtain from (11) $\sigma_{xx}=\sigma_{yy}=0$ and $\sigma_{xy}=\sigma_{yx}=\sigma$ with
$\begin{split}&\sigma=-\frac{\gamma k_{\textsc{b}}}{2}\tau_{\text{rlx}}\omega_{0}\beta^{\prime 2}\iint\limits_{-\pi}^{\;\;\pi}\frac{\mathrm{d}q_{x}\mathrm{d}q_{y}}{(2\pi)^{2}}\\ &\frac{(F_{\bm{q}}+\Omega_{\bm{q}}^{-})(\partial_{q_{x}}F_{\bm{q}}+\partial_{q_{x}}\Omega_{\bm{q}}^{-})(\partial_{q_{y}}F_{\bm{q}}+\partial_{q_{y}}\Omega_{\bm{q}}^{-})}{\sinh^{2}[\frac{\beta^{\prime}}{2}(F_{\bm{q}}+\Omega_{\bm{q}}^{-})]}.\end{split}$ (29)
Rotation of the reference frame by $\pi/4$ transforms the conductivity tensor $\sigma_{\alpha\beta}$ to (12). In the low-temperature limit for the case $\epsilon\ll 1$ we estimate
$\sigma\approx\gamma k_{\textsc{b}}\frac{\omega_{0}\tau_{\text{rlx}}}{\pi}\Delta^{2}\epsilon\mathfrak{c}\ln\frac{1}{1-e^{-\beta^{\prime}\Delta}}.$ (30)
Approximation (30) coincides with (13) up to the multiplier $\mathfrak{c}=-4/[1+\upsilon(2+\kappa)]$.
Appendix D Magnetization for low temperature
Based on the identity $1/(e^{x}-1)=\sum_{n=1}^{\infty}e^{-nx}$, we present $\sinh(\beta^{\prime}\Omega_{q}^{-})/\left[\cosh(\beta^{\prime}F_{q})-\sinh(\beta^{\prime}\Omega_{q}^{-})\right]=1/[e^{\beta^{\prime}(F_{q}-\Omega_{q}^{-})}-1]-1/[e^{\beta^{\prime}(F_{q}+\Omega_{q}^{-})}-1]=\sum_{n=1}^{\infty}[e^{-n\beta^{\prime}(F_{q}-\Omega_{q}^{-})}-e^{-n\beta^{\prime}(F_{q}+\Omega_{q}^{-})}]$. In the low-temperature limit ($\beta^{\prime}/\Delta\gg 1$), the above expression is exponentially localized in the vicinity of $\Gamma$-point. Thus, one can extend the integration over $q_{\alpha}$ in (5) to infinity. Introducing notations $q_{x}=q\cos\chi$ and $q_{y}=q\sin\chi$, we approximate $\Omega_{q}^{-}\approx a(\chi)q^{2}$, $F_{q}\approx\Delta+b(\chi)q^{2}$, where $a(\chi)=(\epsilon_{x}\cos^{2}\chi-\epsilon_{y}\sin^{2}\chi)/8$, and $b(\chi)=[1+(1+\frac{\kappa}{2})(\epsilon_{x}\cos^{2}\chi+\epsilon_{y}\sin^{2}\chi)]/(8\Delta)$. Now one can easily integrate (5) and obtain $\mathcal{M}\approx-\frac{g\mu_{\textsc{b}}}{a_{0}^{2}}\frac{1}{4\pi^{2}\beta^{\prime}}\ln(1-e^{-\beta^{\prime}\Delta})\int_{0}^{2\pi}a(\chi)/[b^{2}(\chi)-a^{2}(\chi)]\mathrm{d}\chi$, where we utilize the identity $\sum_{n=1}^{\infty}e^{-nx}/n=-\ln(1-e^{-x})$. Introducing $\delta\epsilon=(\epsilon_{x}-\epsilon_{y})/2$ and $\bar{\epsilon}=(\epsilon_{x}+\epsilon_{y})/2$, we writhe the piezomagnetic coefficient in form (6) with
$f$ $=\frac{1}{\pi\delta\epsilon}\int\limits_{0}^{\pi}\biggl{\{}\left[1+\left(1+\frac{\kappa}{2}\right)(\bar{\epsilon}+\delta\epsilon\cos\chi)\right]^{2}$ (31a)
$-\Delta^{2}(\delta\epsilon+\bar{\epsilon}\cos\chi)^{2}\biggr{\}}^{-1}(\delta\epsilon+\bar{\epsilon}\cos\chi)\,\mathrm{d}\chi$
In the limit $\delta\epsilon\ll 1$, we obtain
$f=\frac{1}{\left\{\left[1+(1+\frac{\kappa}{2})\bar{\epsilon}\right]^{2}-\Delta^{2}\bar{\epsilon}^{2}\right\}^{3/2}}+\mathcal{O}(\delta\epsilon^{2}).$ (32)
For $\bar{\epsilon}\ll 1$, one approximates $f\approx 1-3(1+\frac{\kappa}{2})\bar{\epsilon}$.